OpenStudy (anonymous):

Find the point on the line 5 x + 4 y + 2 =0 which is closest to the point ( 1, 2 ).

OpenStudy (haseeb96):

4y=-2-5x y=-1\2 -5x\4...........eq #1 m=-5\4 for parallel slope remains same (x,y)=(1,2) put it in the eq#1 2=b -5\4 2=4b-5\4 8=4b-5 8+5=4b 13\4=b b=13\4 y=-5\4 x +13\4